# Law of Large Numbers for continuous $N$-particle ensembles at fixed temperature

### v1

@@ 1.1 Overview
In a series of papers around the year 1990, Voiculescu laid the foundations of what would be known as the theory of Free Probability, from the viewpoint of operator algebras; see e.g. [Voi86]. The subject soon became of main interest in Probability Theory, after unexpected connections with random matrices were discovered, such as the phenomenon of asymptotic freeness [Voi91]. This principle roughly states that large independent Hermitian random matrices behave like free random variables. In particular, the empirical spectral distribution of sums of independent Hermitian random matrices tends to the free convolution of measures, as the matrix dimensions tend to infinity.
@@
By restricting our attention to the spectra of Hermitian matrices, the operation of adding two matrices admits a deformation with the inverse temperature parameter $\beta>0$, in such a way that $\beta=1,2,4$ corresponds to adding self-adjoint real, complex, and quaternionic matrices, respectively. In this paper, we employ the parameter $\theta=\beta/2$, which is more suitable from the point of view of multivariate special functions. The operation is called $\theta$-addition of matrices, though in fact it is an operation on sequences of real numbers (the eigenvalues). One can similarly define $\theta$-projection and $\theta$-multiplication, as the deformations of the operations of cutting corners and multiplying matrices, respectively (see [GM20] for the precise definitions of these three operations). These deformations are natural from the point of view of Statistical Mechanics, which often places particle systems from random matrix eigenvalues as mutually repellent particle systems, e.g. $\beta$-ensembles and log-gas systems. We refer the interested reader to [For10] for a standard reference on these ideas.
@@
The Law of Large Numbers (LLN) of $N$-particle systems, which characterizes the asymptotic global behavior of the empirical measures, has been a central topic in random matrix theory and integrable probability for many years. The idea of using multivariate generating functions of representation theoretic origin on these problems dates back at least to [BuG15], which considers discrete $N$-particles using Schur generating functions. This was later generalized to Jack generating functions with general parameter $\theta>0$ in [Hua21]. In these articles, it was shown that the limits of empirical measures for $\theta$-additions of matrices can be described by the operation of quantized free convolution, instead of free convolution, in the regime of fixed temperature. In recent literature, the high-temperature limit regime, where $\theta\to 0$ at the same time as $N\rightarrow\infty$, so that $N\theta\rightarrow\gamma>0$, has also been a significant source of interest. This setting was studied in [CD25a, CD25b] and [BCG22] for the discrete and continuous settings, respectively.
@@
In this paper, we prove the LLN for the $\theta$-sum of matrices, when $\theta>0$ is fixed and the number $N$ of eigenvalues (the matrix dimension) tends to infinity. We also prove the LLN for the $\theta$-projection of matrices, in the same limit regime. In fact, in our main theorem, we find necessary and sufficient conditions for the LLN for sequences of $N$-tuples of growing size, in terms of the corresponding Bessel generating functions. Our main result answers an open problem posed in [BCG22]; the results for $\theta$-sums and $\theta$-corners are simply applications of the main theorem. We also consider the $\theta$-Dyson Brownian motion started at arbitrary initial conditions.
@@
It is worth noting that our main “if and only if LLN theorem” has been independently proven by Yao [Yao25], but his proof is very different in the “only if direction”. In fact, a new ingredient in our proof that the LLN implies an analytic condition on the sequence of Bessel generating functions of the measures is the recently-discovered topological expansion found by Chapuy–Dolega [CD22].
@@ 1.2 Main result
For the purposes of stating the main result in this introduction, we do not need any exact formulas for the MBF. We simply point out that when $N=1$, the MBF is $B_{(a)}(x;\theta)=e^{ax}$, and when $\theta\to 0$, the MBF becomes
[…]
where $\vec{x}:=(x_{1},\dots,x_{N})$. Also, for $\theta>0$ and the special case when $a_{1}<\dots<a_{N}$, the MBF admits an $\frac{N(N-1)}{2}$-dimensional integral representation [Cue21, Prop. 3.3].
Then to any probability measure $\mu_{N}$ on the closed Weyl chamber $\overline{\mathcal{W}_{N}}=\{(a_{1},\dots,a_{N})\in\mathbb{R}^{N}\mid a_{1}\leq\dots\leq a_{N}\}$, satisfying the technical condition (9) below, we associate the analytic function $G_{\theta}(\vec{x};\mu_{N})$ in a neighborhood of $\big(0^{N}\big)\in\mathbb{C}^{N}$, defined by
@@
This is called the Bessel generating function of $\mu_{N}$. And in view of equation (1), it should be regarded as a symmetric one-parameter $\theta$-deformation of the Fourier transform.
@@
    If this occurs, we shall write: $\lim_{N\rightarrow\infty}{a(N)}\stackrel{{\scriptstyle m}}{{=}}\{m_{k}\}_{k\geq 1}$.
  • Let $a_{\lambda}^{N}$ be the coefficients of $p_{\lambda}(\vec{x})$ in the expansion of $\ln\big(G_{\theta}(\vec{x};\mu_{N})\big)$ in the basis of power sum symmetric functions, for all partitions $\lambda$ with $\ell(\lambda)\leq N$. Then there exist real numbers $\kappa_{1},\kappa_{2},\dots$ such that
@@
      $\lim_{N\to\infty}{\frac{a_{\lambda}^{N}}{N^{\ell(\lambda)}}\bigg|_{\vec{x}=(0^{N})}}=0$, for all partitions $\lambda$ with $\ell(\lambda)\geq 2$.
@@
An important case of application for our theorem is when $\{m_{k}\}_{k\geq 1}$ is the sequence of moments of a unique probability measure $\nu$ on $\mathbb{R}$. Then the convergence (2) implies that the empirical measures $\frac{1}{N}\sum_{i=1}^{N}{\delta_{a_{i}(N)/N}}$ converge weakly, in probability, to $\nu$.
@@ 1.3 Applications: LLN for $\theta$-sums and $\theta$-corners
We briefly outline here the applications of Theorem 3.4, discussed in more detail in Section 6.
Given two independent random $N\times N$ Hermitian matrices $A$, $B$ with uniformly random eigenvectors, and eigenvalues $a_{1},\dots,a_{N}$, and $b_{1},\dots,b_{N}$, respectively, the behavior of the (random) eigenvalues of the sum
@@
as $N\rightarrow\infty$, has been heavily studied in the literature; see e.g. [Voi91], [BG09], [BCG22], and references therein.
For any $\theta>0$, there exists a one-parameter generalization of the above operation on eigenvalues, called the $\theta$-addition; it is defined as follows.
@@
for all $\vec{x}\in\mathbb{R}^{N}$. The generalized function above is compactly supported, so the right side of Eqn. (3) is well-defined. Furthermore, $a+_{\theta}b$ can be equally defined for random $a=(a_{1}\leq\dots\leq a_{N}),b=(b_{1}\leq\dots\leq b_{N})$, by the equality
@@
for all $\vec{x}\in\mathbb{R}^{N}$, whenever the distributions of $a$ and $b$ are two exponentially decaying probability measures (see Definition 2.2), and $\mathbb{E}_{c=a+_{\theta}b}\Big[B_{(c_{1},\dots,c_{N})}(\vec{x};\theta)\Big]$ is again taken by testing the generalized function corresponding to $c$ against the multivariate Bessel function.
Moreover, the distribution of $c=a+_{\theta}b$ is completely determined by the above identity.
It is a well-known conjecture that $a+_{\theta}b$ is not only a generalized function on the Weyl chamber $\mathcal{W}_{N}$, but rather a probability measure; see e.g. [Ank17] for more details. At the discrete level, this positivity conjecture is related to Stanley’s conjecture of positivity of Jack–Littlewood–Richardson coefficients; see [Sta89]. When $\theta=\frac{1}{2},1,2$, equation (4) simply recovers the usual additions of real/complex/quaternionic Hermitian random matrices.
It is a classical result of Voiculescu [Voi91] that, for $\theta=\frac{1}{2},1$, assuming that the empirical spectral measures of the independent $N\times N$ matrices $A(N)$ and $B(N)$ are weakly converging to some deterministic measure $\mu_{a}$ and $\mu_{b}$, respectively, then the empirical measures of $C(N)=A(N)+B(N)$ are also converging to some probability measure $\mu_{c}$. Moreover, this probability measure is
@@
where $\boxplus$ stands for the free convolution, defined through the notion of free cumulants. We refer the reader to [MS17], [NS06] for standard references.
Our first application of Theorem 3.4 is to generalize Voiculescu’s result to $\theta$-additions, for all $\theta>0$. Due to the analytic subtlety of $a+_{\theta}b$, we state our result in terms of moments, i.e, testing the empirical spectral measures with polynomials instead of bounded continuous functions.
@@
Besides addition, another natural operation on an $N\times N$ Hermitian matrix $A$ with eigenvalues $(a_{1},\dots,a_{N})=:a^{(N)}$ is the projection on corners, i.e., the operation of taking the first $M$ rows and columns of $A$ (for some $1\leq M<N$) and forming the $M\times M$ Hermitian sub-matrix $A^{(M)}$. When $A$ has uniformly random eigenvectors, this projection again gives a $M$-tuples of random eigenvalues $(a^{(M)}_{1},...,a^{(M)}_{M})\in\mathbb{R}^{M}$. We can do the same projection on corners for real symmetric and quaternionic Hermitian matrices. In all cases, there is a known explicit joint density function of the eigenvalues $\{a^{(M)}_{i}\}_{1\leq i\leq M\leq N}$ of $A^{(M)}$, where $1\leq M\leq N$, conditioned on $(a^{(N)}_{1}<\dots<a^{(N)}_{N})=(a_{1}<\dots<a_{N})$, where $a_{1},\dots,a_{N}$ are the deterministic eigenvalues of $A$. It turns out that this formula depends on a positive parameter $\theta$, which when specialized to $\theta=\frac{1}{2},1,2$ leads to the cases of real symmetric, complex Hermitian and quaternionic Hermitian matrices, respectively.
A key feature of the corners process just described is that $a^{(k)}\prec a^{(k+1)}$, for all k=1,2,,N1k=1,2,...,N-1, i.e.,
$a^{(k+1)}_{1}\leq a^{(k)}_{1}\leq a^{(k+1)}_{2}\leq\dots\leq a^{(k+1)}_{k}\leq a^{(k)}_{k}\leq a^{(k+1)}_{k+1}$.
So the corners process is actually a probability measure on the interlacing array $a^{(1)}\prec\cdots\prec a^{(N-1)}\prec a^{(N)}=a$ (note that $a^{(N)}=a$ is deterministic). For a general $\theta>0$ and a fixed $a$, one can easily extrapolate the density function from the cases $\theta=\frac{1}{2},1,2$, and get the so-called $\theta$-corners process, as the probability measure on the $\frac{N(N-1)}{2}$ interlacing coordinates $a^{(1)}\prec\cdots\prec a^{(N)}=a$, and with density function
$\begin{split}\Lambda_{\theta}(\vec{y};a)=\frac{1}{Z_{N,\beta,a}}\prod_{k=1}^{N-1}\prod_{1\leq i<j\leq k}|y^{k}_{i}-y_{j}^{k}|^{2-2\theta}\prod_{p=1}^{k}\prod_{q=1}^{k+1}|y_{p}^{k}-y_{q}^{k+1}|^{\theta-1}\cdot\mathbf{1}_{y^{1}\prec y^{2}\prec\cdots\prec y^{N-1}\prec y^{N}=a},\end{split}$ (5)
where $\vec{y}$ is the $\frac{N(N-1)}{2}$-dimensional vector with coordinates $\{y^{k}_{i}\colon 1\leq i\leq k\leq N-1\}$, and $Z_{N,\theta,a}$ is the normalization constant; see e.g. [Cue21, Sec. 2]. This allows us to define the $\theta$-corners as follows.
@@
For $1\leq M<N$, $\theta>0$, and $a=(a_{1}<\dots<a_{N})$ a random real $N$-tuple, the $\theta$-corners of $a$ to the $M^{th}$ level, denoted by $\text{corner}^{N}_{M}(a)$, is the random $M$-tuple whose distribution is the marginal of $y^{(M)}$ in $\Lambda_{\theta}(\vec{y};a)$.
For a (random) $N$-tuple $a(N)=(a_{1}(N)<\dots<a_{N}(N))$, denote its empirical measure by
@@
where the right hand side is the free $\alpha$-projection of $\mu_{a}$, which is the unique probability measure with free cumulants

### v2

@@ 1.1 Overview
In a series of papers around the year 1990, Voiculescu laid the foundations of what would be known as the theory of free probability, from the viewpoint of operator algebras; see e.g. [Voi86]. The subject soon became of main interest in probability theory, after unexpected connections with random matrix theory were discovered, such as the phenomenon of asymptotic freeness [Voi91]. This principle roughly states that large independent Hermitian random matrices behave like free random variables. In particular, the empirical spectral distribution of sums of independent Hermitian random matrices tends to the free convolution of measures, as the matrix dimensions tend to infinity.
@@
By restricting our attention to the spectra of Hermitian matrices, the operation of adding two matrices admits a deformation depending on the real parameter $\beta>0$, in such a way that $\beta=1,2,4$ correspond to adding self-adjoint real, complex, and quaternionic matrices, respectively. In this paper, however, we will employ the deformation parameter $\theta=\beta/2$, which is more suitable from the point of view of multivariate special functions. The resulting operation is then called $\theta$-addition of matrices, though in fact it is an operation on sequences of real numbers (the eigenvalues). One can similarly define $\theta$-projection and $\theta$-multiplication, as the deformations of the operations of cutting corners and multiplying matrices, respectively (see [GM20] for the precise definitions of these three operations). These deformations are natural from the point of view of statistical mechanics, which treats random matrix eigenvalues as mutually repellent random particles and $\beta$ as the inverse temperature, e.g. $\beta$-ensembles and log-gas systems. We refer the interested reader to [For10] for a standard reference on these ideas.
@@
The Law of Large Numbers (LLN) of $N$-particle systems that characterizes the asymptotic global behavior of the empirical measures has been a central topic in random matrix theory and integrable probability for many years. The idea of using multivariate generating functions of representation theoretic origin on these problems dates back at least to [BuG15], which studies discrete $N$-particle systems with the help of Schur generating functions. This was later generalized to Jack generating functions with general parameter $\theta>0$ in [Hua21]. In these articles, it was shown that the limits of empirical measures for $\theta$-additions of matrices can be described by the operation of quantized free convolution, instead of free convolution, in the regime of fixed temperature. In more recent literature, the high-temperature limit regime, where $\theta\to 0$ at the same time as $N\rightarrow\infty$, so that $N\theta\rightarrow\gamma>0$, has also been a significant source of interest. This double scaling regime was studied, for example, in [CD25a, CD25b, CDM] and [BCG22] for discrete and continuous settings, respectively.
@@
In this paper, we prove the LLN for the $\theta$-sum of matrices, in the regime when $\theta>0$ is fixed and the number $N$ of eigenvalues (the matrix dimension) tends to infinity. We also prove the LLN for the $\theta$-projection of matrices, in the same limit regime. In fact, in our main theorem, we find necessary and sufficient conditions for the LLN of sequences of $N$-tuples in terms of the corresponding Bessel generating functions; this result answers an open problem posed by Benaych-Georges, Cuenca and Gorin in [BCG22]. We will then see that the results for $\theta$-sums and $\theta$-corners are simple corollaries. We will also consider the $\theta$-Dyson Brownian motion started at arbitrary initial conditions.
@@
We point out that our main if and only if LLN theorem (namely, Theorem 3.4) has been independently proven by Yao [Yao25]. However, his proof is completely different in one of the directions. In fact, a new ingredient in our proof that the LLN implies an analytic condition on the sequence of Bessel generating functions of the measures is the recently-discovered topological expansion of Chapuy–Dolega [CD22]. This relates the combinatorial objects known as infinite constellations to noncrossing set partitions, or Łukasiewicz paths.
@@ 1.2 Main result
For the purposes of stating the main result in this introduction, we do not need an exact formula for the MBF. We simply point out that if $N=1$, the MBF is $B_{(a)}(x;\theta)=e^{ax}$, while if $\theta\to 0$, the MBF becomes
$\lim_{\theta\to 0}{B_{(a_{1},\dots,a_{N})}(\vec{x};\theta)}=\frac{1}{N!}\sum_{\sigma\in S_{N}}{e^{a_{1}x_{\sigma(1)}+a_{2}x_{\sigma(2)}+\,\dots\,+a_{N}x_{\sigma(N)}}},$ (1)
where $\vec{x}:=(x_{1},\dots,x_{N})$. Also, for $\theta>0$ and the special case when $a_{1}<\dots<a_{N}$, the MBF admits an $\frac{N(N-1)}{2}$-dimensional integral representation; see [Cue21, Prop. 3.3].
Then, for any probability measure $\mu_{N}$ on the closed Weyl chamber $\overline{\mathcal{W}_{N}}=\{(a_{1},\dots,a_{N})\in\mathbb{R}^{N}\mid a_{1}\leq\dots\leq a_{N}\}$ satisfying the technical condition (9) below, we will associate the following analytic function $G_{\theta}(\vec{x};\mu_{N})$ in a neighborhood of $\big(0^{N}\big)\in\mathbb{C}^{N}$:
@@
This is called the Bessel generating function of $\mu_{N}$. In view of equation (1), this should be regarded as a symmetric one-parameter $\theta$-deformation of the Fourier transform of $\mu_{N}$.
@@
    If this occurs, we shall write:
    $\lim_{N\rightarrow\infty}{a(N)}\stackrel{{\scriptstyle m}}{{=}}\{m_{k}\}_{k\geq 1}.$
  • For any partition $\lambda$ with $\ell(\lambda)\leq N$, let $a_{\lambda}^{N}$ be the coefficient of $p_{\lambda}(\vec{x})$ in the expansion of $\ln\big(G_{\theta}(\vec{x};\mu_{N})\big)$ in the basis of power sum symmetric functions. Then there exist real numbers $\kappa_{1},\kappa_{2},\dots$ such that
@@
      $\lim_{N\to\infty}{\frac{a_{\lambda}^{N}}{N^{\ell(\lambda)}}\bigg|_{\vec{x}=(0^{N})}}=0$, whenever $\ell(\lambda)\geq 2$.
@@
An important case of our theorem is when $\{m_{k}\}_{k\geq 1}$ is the sequence of moments of a unique probability measure $\nu$ on $\mathbb{R}$. Then the convergence in Eqn. (2) implies that the empirical measures $\frac{1}{N}\sum_{i=1}^{N}{\delta_{a_{i}(N)/N}}$ converge weakly, in probability, to $\nu$.
@@ 1.3 Applications: LLN for $\theta$-sums and $\theta$-corners
We briefly outline here the applications of Theorem 3.4, to be discussed in more detail in Section 6.
Given two independent random $N\times N$ Hermitian matrices $A$, $B$ with uniformly random eigenvectors, and deterministic eigenvalues $a_{1},\dots,a_{N}$ and $b_{1},\dots,b_{N}$, respectively, the behavior of the (random) eigenvalues of the sum
@@
as $N\rightarrow\infty$, has been heavily studied in the literature; see e.g. [Voi91, BG09, BCG22], and references therein. For a general $\theta>0$, the one-parameter generalization of the above operation on eigenvalues, known as the $\theta$-addition, is defined as follows.
@@
for all $\vec{x}\in\mathbb{R}^{N}$. The generalized function above is compactly supported, so the right side of Eqn. (3) is well-defined. Furthermore, $a+_{\theta}b$ can be equally defined for random $N$-tuples $a=(a_{1}\leq\dots\leq a_{N})$, $b=(b_{1}\leq\dots\leq b_{N})$, by the equality
@@
for all $\vec{x}\in\mathbb{R}^{N}$, whenever the distributions of $a$ and $b$ are exponentially decaying probability measures (see Definition 2.2), and $\mathbb{E}_{c=a+_{\theta}b}\Big[B_{(c_{1},\dots,c_{N})}(\vec{x};\theta)\Big]$ is again taken by testing the generalized function corresponding to $c$ against the multivariate Bessel function.
Moreover, the distribution of $c=a+_{\theta}b$ is uniquely determined by the equality above.
It is a well-known conjecture that $a+_{\theta}b$ is not only a generalized function, but a probability measure; see e.g. [Ank17] for more details. At the discrete level, this positivity conjecture is related to Stanley’s conjecture of positivity of Jack–Littlewood–Richardson coefficients; see [Sta89]. When $\theta=\frac{1}{2},1,2$, equation (4) recovers the usual additions of real, complex and quaternionic Hermitian random matrices.
It is a classical result of Voiculescu [Voi91] that, if $\theta=\frac{1}{2},1$, and assuming that the empirical spectral measures of the independent $N\times N$ matrices $A(N)$ and $B(N)$ weakly converge to some probability measures $\mu_{a}$ and $\mu_{b}$, then the empirical measures of $C(N)=A(N)+B(N)$ also converge to some probability measure $\mu_{c}$. Moreover, this probability measure is the free convolution
@@
that can be defined, for example, through the notion of free cumulants. We refer the reader to [MS17, NS06] for standard references.
Our first application of Theorem 3.4 is the following generalization of Voiculescu’s result, for all $\theta>0$. Due to the analytic subtlety of $a+_{\theta}b$, we state our result in terms of moments, i.e, testing the empirical spectral measures with polynomials instead of bounded continuous functions.
@@
Besides addition, another natural operation on an $N\times N$ Hermitian matrix $A$ with eigenvalues $(a_{1}\leq\dots\leq a_{N})=:a^{(N)}$ is the projection on corners, i.e., the operation of taking the $M\times M$ principal submatrix $A^{(M)}$ consisting of the first $M$ rows and columns of $A$, for some $1\leq M<N$. We will consider the case where the eigenvalues of $A$ are deterministic and the eigenvectors are uniformly random as before; then the random $M$-tuple of real eigenvalues of $A^{(M)}$ will be denoted by $\big(a^{(M)}_{1}\leq...\leq a^{(M)}_{M}\big)=:a^{(M)}$. We can do the same projection on corners for real symmetric and quaternionic Hermitian matrices. In all cases, by considering all principal submatrices $A^{(M)}$, as $M$ ranges over $1,2,\dots,N$, we obtain a triangular array $\big\{a^{(M)}_{i}\colon 1\leq i\leq M\leq N\big\}$ of random eigenvalues, where the top row $a^{(N)}$ is the fixed set of eigenvalues of $A$. This random array is known as the corners process, and a key feature of it is the interlacing condition
$a^{(k+1)}_{1}\leq a^{(k)}_{1}\leq a^{(k+1)}_{2}\leq\dots\leq a^{(k+1)}_{k}\leq a^{(k)}_{k}\leq a^{(k+1)}_{k+1}$,
denoted by $a^{(k)}\prec a^{(k+1)}$, and valid for all k=1,,N1k=1,...,N-1.
In all cases, the joint density function for the array of eigenvalues $\big\{a^{(M)}_{i}\big\}$ can be calculated explicitly. It turns out that this formula can be naturally expressed as a function of a positive parameter $\theta$, such that for $\theta=\frac{1}{2},1,2$, it specializes to the cases of real symmetric, complex Hermitian and quaternionic Hermitian matrices, respectively. Moreover, for a general $\theta>0$, this formula is the density of an actual probability measure on the set of interlacing arrays $a^{(1)}\prec\cdots\prec a^{(N-1)}\prec a^{(N)}$, where $a^{(N)}=a$ is deterministic. This measure is called the $\theta$-corners process; directly, it can be defined as the probability measure on the set of $\frac{N(N-1)}{2}$ interlacing coordinates $a^{(1)}\prec\cdots\prec a^{(N)}=a$, with density function
$\Lambda_{\theta}(\vec{y};a):=\frac{1}{Z_{N,\theta,a}}\prod_{k=1}^{N-1}\prod_{1\leq i<j\leq k}|y^{k}_{i}-y_{j}^{k}|^{2-2\theta}\prod_{p=1}^{k}\prod_{q=1}^{k+1}|y_{p}^{k}-y_{q}^{k+1}|^{\theta-1}\cdot\mathbf{1}_{y^{1}\prec y^{2}\prec\cdots\prec y^{N-1}\prec y^{N}=a},$ (5)
where $\vec{y}$ is the $\frac{N(N-1)}{2}$-dimensional vector with coordinates $\{y^{k}_{i}\colon 1\leq i\leq k\leq N-1\}$, and where $Z_{N,\theta,a}$ is the normalization constant; see e.g. [Cue21, Sec. 2]. We will assume the strict inequalities $a_{1}<\dots<a_{N}$, so that no particle $y^{k}_{i}$ is deterministic.
@@
For $1\leq M<N$, $\theta>0$, and $a=(a_{1}<\dots<a_{N})$ a random real $N$-tuple, the random $M$-tuple whose distribution is the marginal of $y^{(M)}$, according to the distribution $\Lambda_{\theta}(\vec{y};a)$, is denoted by $\text{corner}^{N}_{M}(a)$ and called the $M^{th}$ level marginal of the $\theta$-corners process of $a$.
For a random $N$-tuple $a(N)=(a_{1}(N)<\dots<a_{N}(N))$, we denote its empirical measure by
@@
where the right hand side is the free $\alpha$-projection of $\mu_{a}$, defined as the unique probability measure with free cumulants